The supervisor of a Muesli Bar producing factory noted that the weight of each 30g bar is

a normally distributed random variable with a mean of 30.25g and a standard deviation of

0.2g. Find the probability that the mean weight of 5 Muesli bars is less than 30g. Suppose

the distribution of weight of each Muesli bar is non-normal. Under what situation would

the above answer not, change?


1
Expert's answer
2021-11-22T13:53:17-0500

a) Let "X=" the mean weight of "n" Muesli bars: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=30.25\ g, \sigma=0.2\ g, n=5"


"P(X<30)=P(Z<\dfrac{30-30.25}{0.2/\sqrt{5}})"

"\approx P(Z<-2.795085)\approx0.0025943"

b) By the Central Limit Theorem the normal approximation for "X" will generally be good if "n \geq 30." If "n < 30," the approximation is good only if the population is not too different from a normal distribution (if the population is known to be normal, the sampling distribution of "X" will follow a normal distribution exactly, no matter how small the size of the samples.

The sampling distribution of "X" will still be approximately normal with mean "\mu_X=30" and variance "s^2\/n=0.2^2\/5," provided that the sample size "n" is large "(n \geq 30)."


"P(X<30)=P(Z<\dfrac{30-30.25}{s/\sqrt{n}})"

"=P(Z<\dfrac{30-30.25}{0.2/\sqrt{5}})\approx0.0025943"